OpenStudy (anonymous):

if x equals -3, then find the matching value for y. -7x-2y=3

OpenStudy (bahrom7893):

Okay for this one. Plug in x=-3. What do u get?

OpenStudy (bahrom7893):

-7(-3)-2y=3

OpenStudy (bahrom7893):

Now simplify that and post what it looks like.

OpenStudy (anonymous):

i got a weird answer?

OpenStudy (anonymous):

wut did u get?

OpenStudy (bahrom7893):

okay wait up.. no i get a nice answer.

OpenStudy (bahrom7893):

Post the steps here ill showu the mistakes

OpenStudy (anonymous):

14

OpenStudy (bahrom7893):

jlear first of all don't just post the answer

OpenStudy (bahrom7893):

2nd of all post the right answer the answer is wrong

OpenStudy (bahrom7893):

And jlear finally if u see someone who knows what he/she is doing just move on to the next question.

OpenStudy (bahrom7893):

the answer is 9

OpenStudy (bahrom7893):

I could've given that in 10 seconds, but I want them to learn.

OpenStudy (bahrom7893):

So even posting the solution doesn't help. Give them prompts and let them work.

OpenStudy (anonymous):

ohhh, sorry, I'm new lol... its not 1/14? and will do.

OpenStudy (anonymous):

sorry nvm i just read your equation, just now read the question, but will do from now on, sry bahrom

OpenStudy (bahrom7893):

Okay here: -7(-3)-2y=3 21-2y=3 -2y=3-21 -2y=-18 y=-18/-2 y=9

OpenStudy (anonymous):

yeh ur right xD